# The Uniform Mordell–Lang Conjecture

### v3

@@ 5.1. A result in family
Before moving on, let us take a closer look at the constants obtained in the first three steps. First of all, the two integers $m=m(S)>0$ and $M=M(S)>0$ were taken in Step 1 and depend only on $S$. We obtain in Step 3 a proper closed subvariety $Z_{S^{\prime}}$ of $\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ and a constant $\delta_{S^{\prime}}$ for each irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$; à priori they depend on the function $f_{S^{\prime},1}$ and the constant $\epsilon_{S^{\prime}}>0$ chosen in Step 2. However, if $S^{\prime}$, $m$ and $M$ are fixed, then the measures $\mu_{S^{\prime},1}$ and $\mu_{S^{\prime},2}$ in Step 2 are fixed, and so one can fix a choice of $f_{S^{\prime},1}$ and $\epsilon_{S^{\prime}}>0$. Therefore, we can and do view $Z_{S^{\prime}}$ and $\delta_{S^{\prime}}>0$ to depend only on $S^{\prime}$, $m$ and $M$.
Step 4 Conclude with a similar argument for Proposition 4.2.
Again, we are divided into two cases.
Consider the abelian scheme $\pi^{[m(M+2)]}\colon\mathcal{A}_{\mathbf{H}^{\circ}}^{[m(M+2)]}\rightarrow\mathbf{H}^{\circ}$.
Consider $S^{\prime}\setminus\pi^{[m(M+2)]}(\mathscr{X}_{S^{\prime}}^{[m(M+2)]}\setminus Z_{S^{\prime}})$ endowed with the reduced induced subscheme structure; it has dimension $\leq\dim S^{\prime}-1\leq\dim S-1$. Let $S_{1},\ldots,S_{k}$ be the irreducible components of $\bigcup_{S^{\prime}}S^{\prime}\setminus\pi^{[m(M+2)]}(\mathscr{X}_{S^{\prime}}^{[m(M+2)]}\setminus Z_{S^{\prime}})$ with $S^{\prime}$ running over all irreducible components of $\overline{S_{\mathrm{gen}}}$. Then the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$ and $m$ and $M$. But $m=m(S)>0$ and $M=M(S)>0$ depend only on $S$. So the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$.
Case (i) $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$.
For each $i\in\{1,\ldots,k\}$, we can apply the induction hypothesis to $S_{i}$ since $\dim S_{i}\leq\dim S-1$. So we obtain constants $c_{i,2}^{\prime\prime}=c_{i,2}^{\prime\prime}(r,d,S_{i})>0$ and $c_{i,3}^{\prime\prime}=c_{i,3}^{\prime\prime}(r,d,S_{i})>0$ such that for each $s\in S_{i}(\overline{\mathbb{Q}})$, the set
@@
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{i,2}^{\prime\prime}$.
Let $c_{\mathrm{deg},2}^{\prime\prime}=\max_{1\leq i\leq k}\{c_{i,2}^{\prime\prime}\}>0$ and $c_{\mathrm{deg},3}^{\prime\prime}=\min_{1\leq i\leq k}\{c_{i,3}^{\prime\prime}\}>0$. We have seen above that the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$. So $c_{\mathrm{deg},2}^{\prime\prime}$ and $c_{\mathrm{deg},3}^{\prime\prime}$ thus obtained depend only on $r$, $d$ and $S$. Moreover, for each $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{i,s}$ for some $i$; so $\Sigma_{\mathrm{deg},s}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{\mathrm{deg},2}^{\prime\prime}$. This concludes for Case (i).
Case (ii) $s\in\pi^{[m(M+2)]}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$ for some irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$. In this case, $(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})_{s}\not=\emptyset$. So $\mathscr{X}_{s}^{m(M+2)}=(\mathscr{X}_{S^{\prime}}^{[m(M+2)]})_{s}\not=(Z_{S^{\prime}})_{s}$.
Set $c_{S^{\prime},3}^{\prime\prime}=\delta_{S^{\prime}}/4m(M+2)$. Then $c_{S^{\prime},3}^{\prime\prime}$ depend only on $S^{\prime}$, $m$ and $M$. Consider
@@
We claim that $\Sigma_{S^{\prime},s}^{m(M+2)}\subseteq(Z_{S^{\prime}})_{s}(\overline{\mathbb{Q}})$. Indeed assume not, then there exists $\mathbf{x}:=(x_{1},\ldots,x_{m(M+2)})\in(\Sigma_{S^{\prime},s}^{m(M+2)}\setminus(Z_{S^{\prime}})_{s})(\overline{\mathbb{Q}})$. Then $\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+2)}}(\mathbf{x})=\sum_{i=1}^{m(M+2)}\hat{h}_{{\mathcal{L}}}(x_{i})\leq m(M+2)c_{S^{\prime},3}^{\prime\prime}<\delta_{S^{\prime}}$. On the other hand, each component of $\mathscr{D}(\mathbf{x})$ is of the form $x_{k}$ or of the form $x_{i}-x_{j}$ for some $i$ and $j$, and $\hat{h}_{{\mathcal{L}}}(x_{j}-x_{i})\leq 2\hat{h}_{{\mathcal{L}}}(x_{j})+2\hat{h}_{{\mathcal{L}}}(x_{i})\leq 4c_{S^{\prime},3}^{\prime\prime}$. So $\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+1)}}(\mathscr{D}(\mathbf{x}))\leq m(M+1)4c_{S^{\prime},3}^{\prime\prime}<\delta_{S^{\prime}}$. Thus we have reached a contradiction to the conclusion of Step 3. This establishes the claim.
Recall the assumption $\mathscr{X}_{s}^{m(M+2)}\not=(Z_{S^{\prime}})_{s}$ for this case. Thus we are allowed to apply Lemma 4.3 to $X=\mathscr{X}_{s}$, $L={\mathcal{L}}_{s}|_{\mathscr{X}_{s}}$, $Z=(Z_{S^{\prime}})_{s}$ and $\Sigma=\Sigma_{S^{\prime},s}$. So there exists a Zariski closed subset $X^{\prime}$ of $\mathscr{X}_{s}$ such that
@@
    $X^{\prime}\subsetneq\mathscr{X}_{s}$;
@@
But $\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}(Z_{S^{\prime}})_{s}$ depends only on $S^{\prime}$, $m$ and $M$. Hence property (ii) above can be simplified to be $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c_{S^{\prime},2}^{\prime\prime}$, with $c_{S^{\prime},2}^{\prime\prime}$ depending only on $r$, $d$, $S^{\prime}$, $m$ and $M$. In summary, the set $\Sigma_{S^{\prime},s}$ defined in (5.10) is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c_{S^{\prime},2}^{\prime\prime}$.
Let $c_{2}^{*\prime\prime}=\max_{S^{\prime}}\{c_{S^{\prime},2}^{\prime\prime}\}>0$ and $c_{3}^{*\prime\prime}=\min_{S^{\prime}}\{c_{S^{\prime},3}^{\prime\prime}\}>0$, where $S^{\prime}$ runs over all irreducible components of $\overline{S_{\mathrm{gen}}}$ defined by (3.5). The subset $S_{\mathrm{gen}}$ of $S$ is uniquely determined by $S$, and hence the set $\{S^{\prime}\}$ is uniquely determined by $S$. Moreover $m=m(S)>0$ and $M=M(S)>0$ depend only on $S$. So the constants $c_{2}^{*\prime\prime}$ and $c_{3}^{*\prime\prime}$ thus obtained depend only on $r$, $d$ and $S$. The discussion above yields the following assertion: for each $s\in\bigcup_{S^{\prime}}\pi^{[m]}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{S^{\prime},s}$ for some $S^{\prime}$; so $\Sigma_{s}^{*}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{*\prime\prime}$. This concludes for Case (ii).
Now let $s\in S_{\mathrm{gen}}$ be an arbitrary point. Then we are either in Case (i) or Case (ii). So the proposition holds true by letting $c_{2}^{\prime\prime}=\max\{c_{\mathrm{deg},2}^{\prime\prime},c_{2}^{*\prime\prime}\}>0$ and $c_{3}^{\prime\prime}=\min\{c_{\mathrm{deg},3}^{\prime\prime},c_{3}^{*\prime\prime}\}>0$; all the constants involved depend only on $r$, $d$ and $S$ (see (5.10) and (5.11)). ∎
@@
Let $(A,L)$ be a polarized abelian varieties defined over $\overline{\mathbb{Q}}$ with $\deg_{L}A=l$. Let $X$ be an irreducible subvariety of $A$, defined over $\overline{\mathbb{Q}}$, with $\dim X=r$ and $\deg_{L}X=d$ such that $X$ generates $A$. We may and do assume $r\geq 1$ because otherwise the proposition is trivial.
We start with the case where $(A,L)$ is principally polarized; notice that $l=g!$.
@@
Now $(A,L)$ gives rise to a point $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$ such that $(\mathfrak{A}_{g})_{b}\cong A$ and $c_{1}(\mathfrak{L}_{g}|_{A})=2c_{1}(L)=c_{1}(L\otimes L_{-})$ for $L_{-}=[-1]^{*}L$. We have $\deg_{\mathfrak{L}_{g}^{\otimes 4}|_{A}}X=8^{r}d$ by (4.10).
@@
${\mathscr{X}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\iota}$${\mathfrak{A}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\mathrm{univ}}}$${\mathbf{H}^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{\mathbf{H}^{\circ}}}$${\mathbb{A}_{g}}$
@@
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. All objects in this diagram are varieties defined over $\overline{\mathbb{Q}}$. Then the pair $(X,(A,L))$ is parametrized by a point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$. Thus $X=\mathscr{X}_{s}$ and $A=\mathcal{A}_{s}=(\mathfrak{A}_{g})_{b}$, and $\iota_{\mathbf{H}^{\circ}}(s)=b$.
For the line bundle ${\mathcal{L}}=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$, we have seen that $\hat{h}_{{\mathcal{L}}|_{A}}=4\hat{h}_{L\otimes L_{-}}$ as height functions on $A(\overline{\mathbb{Q}})$ in (4.12).
@@
Let $\mathbf{H}^{\circ}_{\mathrm{gen}}$ be the subset of $\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ as defined by (3.9). If the stabilize of $X$ in $A$ (denoted by $\mathrm{Stab}_{A}(X)$) has positive dimension, then $X^{\circ}=\emptyset$ by definition of Ueno locus and hence the proposition trivially holds true. Thus we may and do assume that $\mathrm{Stab}_{A}(X)$ is finite. Then the point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ which parametrizes $(X,(A,L))$ is in $\mathbf{H}^{\circ}_{\mathrm{gen}}$.
@@
Apply Proposition 5.2 to $S=\mathbf{H}^{\circ}$. Then we obtain constants $c_{2}^{\prime\prime}$ and $c_{3}^{\prime\prime}$ depending only on $r$, $8^{r}d$ and $\mathbf{H}^{\circ}$ (and hence only on $g$, $d$ and $r$). The set $\Sigma:=\left\{x\in\mathscr{X}_{s}^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{{\mathcal{L}}|_{A}}(x)\leq c_{3}^{\prime\prime}\right\}=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{3}^{\prime\prime}/4\right\}$ is contains in $X^{\prime}$, for some proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{{\mathcal{L}}|_{A}}(X^{\prime})\leq c_{2}^{\prime\prime}$. Thus Proposition 5.1 for this case holds true because $\deg_{L}X^{\prime}<\deg_{{\mathcal{L}}|_{A}}(X^{\prime})$ (since $c_{1}({\mathcal{L}}|_{A})=4c_{1}(\mathfrak{L}_{g}|_{A})=8c_{1}(L)$).
@@
Now let us turn to arbitrary $(A,L)$. By Lemma 2.2.(iv) and the inverse isogeny, there exist a principally polarized abelian variety $(A_{0},L_{0})$ defined over $\overline{\mathbb{Q}}$ and an isogeny $u\colon A_{0}\rightarrow A$ such that $u^{*}L=[l/g!]^{*}L_{0}$; see below (4.13). In particular, $u^{*}(L\otimes L_{-})\cong(L_{0}\otimes(L_{0})_{-})^{\otimes(l/g!)^{2}}$.
@@
Let $X_{0}$ be an irreducible component of $u^{-1}(X)$. Then $\deg_{u^{*}L}X_{0}\leq d(l/g!)^{2g-1}$ by (4.14), and $X_{0}$ is not contained in any proper subgroup of $A_{0}$. Thus we can apply the conclusion for the principally polarized case to $X_{0}\subseteq A_{0}$ and get two constants $c_{2}^{\prime\prime}$ and $c_{3}^{\prime\prime}$ depending only on $g$, $r$, $l$ and $d$.[4][4][4]Again, we first of all get constants $c_{0,2}^{\prime\prime}(g,r,d^{\prime})>0$ and $c_{0,3}^{\prime\prime}(g,r,d^{\prime})>0$ for each $1\leq d^{\prime}\leq d$, and then set $c_{2}^{\prime\prime}:=\max_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,2}^{\prime\prime}(g,r,d^{\prime})\}$ and $c_{3}^{\prime\prime}:=\min_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,3}^{\prime\prime}(g,r,d^{\prime})\}$.
By the definition of Ueno locus, we have $X^{\circ}=u(X_{0})^{\circ}\subseteq u(X_{0}^{\circ})$. So we have $\Sigma\subseteq u(\Sigma_{0})$ for $\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{3}^{\prime\prime}\right\}$ and $\Sigma_{0}=\left\{x_{0}\in X_{0}^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L_{0}\otimes(L_{0})_{-}}(x_{0})\leq(l/g!)^{2}c_{3}^{\prime\prime}\right\}$. Now the conclusion follows from the principally polarized case with $c_{3}^{\prime\prime}$ replaced by $(l/g!)^{2}c_{3}^{\prime\prime}$.
@@
Let $g$, $l$, $r$, and $d$ be positive integers. There exist constants $c_{1}=c_{1}(g,l,r,d)>0$ and $c_{2}=c_{2}(g,l,r,d)>0$ satisfying the following property. For

### v4

@@ 5.1. A preliminary result for the restricted Hilbert scheme
Indeed, assume both conditions were violated. Then (5.8) would hold true, and (5.9) with $\mathbf{x}^{\prime}=\mathscr{D}(\mathbf{x})$ would also hold true.
But
$\frac{1}{\#\mathbf{O}(\mathbf{x})}\sum_{\mathbf{y}\in\mathbf{O}(\mathbf{x})}f_{S^{\prime},1}(\mathbf{y})=\frac{1}{\#\mathbf{O}(\mathscr{D}(\mathbf{x}))}\sum_{\mathbf{y}^{\prime}\in\mathbf{O}(\mathscr{D}(\mathbf{x}))}f_{S^{\prime},2}(\mathbf{y}^{\prime})$
because $f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$. So we would have
$\left|\int_{\mathscr{X}^{[m(M+2)]}_{S^{\prime}}(\mathbb{C})}f_{S^{\prime},1}\mu_{S^{\prime},1}-\int_{\mathscr{D}(\mathscr{X}^{[m(M+2)]}_{S^{\prime}})(\mathbb{C})}f_{S^{\prime},2}\mu_{S^{\prime},2}\right|\leq 2\varepsilon_{S^{\prime}},$
which contradicts (5.7) as $f_{S^{\prime},1}=f_{S^{\prime},2}\circ\mathscr{D}$. This finishes the proof of the claim.
Before moving on, let us take a closer look at the constants and objects obtained up to now. In Step 1, we introduced the integers $m=m(S)>0$ and $M=M(S)>0$, which depend only on $S$. The measures $\mu_{S^{\prime},1}$ and $\mu_{S^{\prime},2}$ from Step 2 clearly depend only on $S^{\prime}$, $m$ and $M$, and so do the test functions $f_{S^{\prime},1}$, $f_{S^{\prime},2}$ and the constant $\varepsilon_{S^{\prime}}>0$ chosen in Step 3. In this step, we obtained a closed subvariety $Z_{S^{\prime}}\subsetneq\mathscr{X}^{[m(M+2)]}_{S^{\prime}}$ and a constant $\delta_{S^{\prime}}$ for each irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$; these depend only on $f_{S^{\prime},1}$, $f_{S^{\prime},2}$ and $\varepsilon_{S^{\prime}}>0$ and hence ultimately only on $S^{\prime}$, $m$ and $M$.
Step 5: Conclusion and induction step
The argument of this step is similar to the proof of Proposition 4.2. The main difference is that the the height inequality [DGH21, Thm. 1.6 and B.1] is replaced by the height lower bound from Step 4.
By abuse of notation, let $\pi\colon\mathcal{A}_{S^{\prime}}\rightarrow S^{\prime}$ denote the structural morphism. Consider the complement $S^{\prime}\setminus\pi(\mathscr{X}_{S^{\prime}}^{[m(M+2)]}\setminus Z_{S^{\prime}})$ endowed with its reduced induced subscheme structure; it has dimension $\leq\dim S^{\prime}-1\leq\dim S-1$. Let $S_{1},\ldots,S_{k}$ be the irreducible components of $\bigcup_{S^{\prime}}S^{\prime}\setminus\pi(\mathscr{X}_{S^{\prime}}^{[m(M+2)]}\setminus Z_{S^{\prime}})$ with $S^{\prime}$ running over all irreducible components of $\overline{S_{\mathrm{gen}}}$. The set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$, $m$ and $M$. As both $m$ and $M$ depend only on $S$, the set $\{S_{1},\ldots,S_{k}\}$ is actually completely determined by $S$.
As in the proof of Proposition 4.2, we need to distinguish two cases in order to prove the proposition for a given $s\in S_{\mathrm{gen}}(\overline{\mathbb{Q}})$.
Case (i): $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$.
Since $\dim S_{i}\leq\dim S-1$ for all $i\in\{1,\ldots,k\}$, we can apply the induction hypothesis for each subvariety $S_{i}$ separately. This yields constants $c_{i,2}^{\prime\prime}=c_{i,2}^{\prime\prime}(r,d,S_{i})>0$ and $c_{i,3}^{\prime\prime}=c_{i,3}^{\prime\prime}(r,d,S_{i})>0$ such that, for each $s\in S_{i}(\overline{\mathbb{Q}})$, the set
@@
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{i,2}^{\prime\prime}$.
Set $c_{\mathrm{deg},2}^{\prime\prime}:=\max_{1\leq i\leq k}\{c_{i,2}^{\prime\prime}\}>0$ and $c_{\mathrm{deg},3}^{\prime\prime}:=\min_{1\leq i\leq k}\{c_{i,3}^{\prime\prime}\}>0$. As the set $\{S_{1},\ldots,S_{k}\}$ is completely determined by $S$, the constants $c_{\mathrm{deg},2}^{\prime\prime}$ and $c_{\mathrm{deg},3}^{\prime\prime}$ depend only on $r$, $d$ and $S$.
For each $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{i,s}$ for some $i$. Therefore, $\Sigma_{\mathrm{deg},s}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{\mathrm{deg},2}^{\prime\prime}$. This proves the assertion of the proposition in this case.
Case (ii): $s\in\pi(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$ for some irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$.
Set $c_{S^{\prime},3}^{\prime\prime}=\delta_{S^{\prime}}/4m(M+2)$ and consider
@@
We claim that
(5.13) $\Sigma_{S^{\prime},s}^{m(M+2)}\subseteq Z_{s}(\overline{\mathbb{Q}}).$
Assume to the contrary that there exists some $\mathbf{x}=(x_{1},\ldots,x_{m(M+2)})\in(\Sigma_{S^{\prime},s}^{m(M+2)}\setminus Z_{s})(\overline{\mathbb{Q}})$. Then,
$\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+2)}}(\mathbf{x})=\sum_{i=1}^{m(M+2)}\hat{h}_{{\mathcal{L}}}(x_{i})\leq m(M+2)c_{S^{\prime},3}^{\prime\prime}<\delta_{S^{\prime}}$
and
$\hat{h}_{{\mathcal{L}}^{\boxtimes m(M+1)}}(\mathscr{D}(\mathbf{x}))=\sum_{i=1}^{m}\hat{h}_{{\mathcal{L}}}(x_{i})+\sum_{i=2}^{M}\sum_{j=1}^{m}\hat{h}_{{\mathcal{L}}}(x_{i\cdot m+j}-x_{m+j})\leq 4m(M+1)c_{S^{\prime},3}^{\prime\prime}<\delta_{S^{\prime}}.$
This contradicts the alternative height bounds obtained in Step 4, establishing the claim.
By the assumption of this case, $\mathscr{X}_{s}^{m(M+2)}\not=Z_{s}$. Therefore, we can apply Lemma 4.3 to $X=\mathscr{X}_{s}$, $L={\mathcal{L}}_{s}|_{\mathscr{X}_{s}}$, $Z=Z_{s}$ and $\Sigma=\Sigma_{S^{\prime},s}$. In this way, we obtain a proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ such that
@@
    $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c(m(M+2),r,d,\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}Z_{s})$, and
@@
As the degree $\deg_{{\mathcal{L}}_{s}^{\boxtimes m}}Z_{s}$ depends only on $S^{\prime}$, $m$ and $M$, property (i) simplifies to $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c_{S^{\prime},2}^{\prime\prime}$ with a constant $c_{S^{\prime},2}^{\prime\prime}$ depending only on $r$, $d$, $S^{\prime}$, $m$ and $M$. In summary, the set $\Sigma_{S^{\prime},s}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for some Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})\leq c_{S^{\prime},2}^{\prime\prime}$.
Set $c_{2}^{*\prime\prime}:=\max_{S^{\prime}}\{c_{S^{\prime},2}^{\prime\prime}\}>0$ and $c_{3}^{*\prime\prime}:=\min_{S^{\prime}}\{c_{S^{\prime},3}^{\prime\prime}\}>0$ where $S^{\prime}$ runs over all irreducible components of $\overline{S_{\mathrm{gen}}}$. The subset $S_{\mathrm{gen}}$ of $S$ is uniquely determined by $S$. Hence the set $\{S^{\prime}\}$ of irreducible components of $\overline{S_{\mathrm{gen}}}$ is uniquely determined by $S$. Moreover $m=m(S)>0$ and $M=M(S)>0$ depend only on $S$. Consequently, the constants $c_{2}^{*\prime\prime}$ and $c_{3}^{*\prime\prime}$ depend only on $r$, $d$ and $S$.
For each $s\in\bigcup_{S^{\prime}}\pi(\mathscr{X}^{[m(M+2)]}_{S^{\prime}}\setminus Z_{S^{\prime}})(\overline{\mathbb{Q}})$, the set
@@
must be contained in $\Sigma_{S^{\prime},s}$ for some $S^{\prime}$. So $\Sigma_{s}^{*}$ is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{*\prime\prime}$. This concludes the induction step in this case. ∎
@@
Let $(A,L)$ be a polarized abelian variety defined over $\overline{\mathbb{Q}}$ with $\deg_{L}A=l$. Let $X$ be an irreducible subvariety of $A$ with $\dim X=r$ and $\deg_{L}X=d$ such that $X$ generates $A$. We may assume $r\geq 1$; otherwise the proposition is trivial.
We start with the case where $(A,L)$ is principally polarized, i.e. $l=g!$.
@@
[…]
@@
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. All varieties and morphisms in this diagram are defined over $\overline{\mathbb{Q}}$. The pair $(X,(A,L))$ is parametrized by a point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$, which means that $X=\mathscr{X}_{s}$, $A=\mathcal{A}_{s}=(\mathfrak{A}_{g})_{b}$, and $\iota_{\mathbf{H}^{\circ}}(s)=b$.
@@
Let $\mathbf{H}^{\circ}_{\mathrm{gen}}$ be the subset of $\mathbf{H}^{\circ}$ as defined by (3.9). If the stabilizer $\mathrm{Stab}_{A}(X)$) of $X$ in $A$ has positive dimension, then $X^{\circ}=\emptyset$ by definition of the Ueno locus and the proposition trivially holds true. Thus we may and do assume that $\mathrm{Stab}_{A}(X)$ is finite. Then the point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ which parametrizes $(X,(A,L))$ is in $\mathbf{H}^{\circ}_{\mathrm{gen}}$
We apply Proposition 5.2 to $S=\mathbf{H}^{\circ}$, and hence obtain constants $c_{2}^{\prime\prime}$ and $c_{3}^{\prime\prime}$ that depend only on $r$, $8^{r}d$ and $\mathbf{H}^{\circ}$ (so only on $g$, $d$ and $r$) such that the set
$\Sigma:=\left\{x\in\mathscr{X}_{s}^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{{\mathcal{L}}|_{A}}(x)\leq c_{3}^{\prime\prime}\right\}=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{3}^{\prime\prime}/4\right\}$
@@
is contained in some proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{{\mathcal{L}}|_{A}}(X^{\prime})\leq c_{2}^{\prime\prime}$. Thus Proposition 5.1 for this case holds true because $\deg_{L}X^{\prime}<\deg_{{\mathcal{L}}|_{A}}(X^{\prime})$ (since $c_{1}({\mathcal{L}}|_{A})=4c_{1}(\mathfrak{L}_{g}|_{A})=8c_{1}(L)$).
@@
Now let us turn to arbitrary polarized abelian varieties $(A,L)$. By Lemma 2.2.(iv), there exist a principally polarized abelian variety $(A_{0},L_{0})$ defined over $\overline{\mathbb{Q}}$ and an isogeny $u\colon A_{0}\rightarrow A$ such that $u^{*}L=[l/g!]^{*}L_{0}$; see below (4.13). In particular, $u^{*}(L\otimes L_{-})\cong(L_{0}\otimes(L_{0})_{-})^{\otimes(l/g!)^{2}}$, where $(L_{0})=[-1]^{*}L_{0}$.
@@
Let $X_{0}$ be an irreducible component of $u^{-1}(X)$. Then $\deg_{u^{*}L}X_{0}\leq d(l/g!)^{2g-1}$ by (4.14), and $X_{0}$ is not contained in any proper subgroup of $A_{0}$. Thus we can apply the conclusion for the principally polarized case to $X_{0}\subseteq A_{0}$ and get two constants $c_{2}^{\prime\prime}$ and $c_{3}^{\prime\prime}$ depending only on $g$, $r$, $l$ and dd.[3][3][3]Again, we first of all get constants $c_{0,2}^{\prime\prime}(g,r,d^{\prime})>0$ and $c_{0,3}^{\prime\prime}(g,r,d^{\prime})>0$ for each $1\leq d^{\prime}\leq d$, and then set $c_{2}^{\prime\prime}:=\max_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,2}^{\prime\prime}(g,r,d^{\prime})\}$ and $c_{3}^{\prime\prime}:=\min_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,3}^{\prime\prime}(g,r,d^{\prime})\}$.
By definition of the Ueno locus, we have $X^{\circ}=u(X_{0})^{\circ}\subseteq u(X_{0}^{\circ})$. So we have $\Sigma\subseteq u(\Sigma_{0})$ for
$\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{3}^{\prime\prime}\right\}$
and
$\Sigma_{0}=\left\{x_{0}\in X_{0}^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L_{0}\otimes(L_{0})_{-}}(x_{0})\leq(l/g!)^{2}c_{3}^{\prime\prime}\right\}.$
The conclusion follows from the principally polarized case upon replacing $c_{3}^{\prime\prime}$ by $(l/g!)^{2}c_{3}^{\prime\prime}$. ∎
@@
Let $g$, $l$, $r$, and $d$ be positive integers. There exist constants $c_{1}=c_{1}(g,l,r,d)>0$ and $c_{2}=c_{2}(g,l,r,d)>0$ with the following property. For